Craig and Tom, working together, can clean the house in 8 hours. Working alone, Tom takes four times as long as Craig. How long does it take Craig to clean the house alone?
Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

Let Craig be c . Then Tom = 4c. So:
1/c+1/4c=1/8
8+2=c
c=10
Craig can clean the house in 10 hours, and Tom can clean the house in 40 hours. ☺☺☺☺

Answer by ikleyn(52908)